# Time Weaver: A Conditional Time Series Generation Model

### v1

@@ 7 Conclusion
This paper addresses a critical gap in synthetic time series data generation by introducing Time Weaver, a novel diffusion-based generative model. Time Weaver leverages heterogeneous paired metadata, encompassing categorical, continuous, and time-variant variables, to significantly improve the quality of generated time series. Moreover, we introduce a new evaluation metric, J-FTSD, to assess conditional time series generation models. This metric offers a refined approach to evaluating the specificity of generated time series relative to paired metadata conditions. Through Time Weaver , we demonstrate state-of-the-art results across four diverse real-world datasets.
@@
Appendix
[…]
Appendix A Diffusion Process
DMs are trained to denoise a noisy sample, referred to as the backward process $p_{\theta}$, generated by a Markovian forward process $q$. The forward process is predetermined by specifying a noise schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$. The following equations parameterize the forward process:
@@ Appendix A Diffusion Process
where $x_{0}\sim\mathcal{X}$ and $T$ is the number of diffusion steps. The noise schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$ and $T$ are chosen such that the distribution of $x_{T}$ is zero-mean, unit-variance normal distribution, i.e., $q(x_{T})\simeq\mathcal{N}(\mathbf{0},\mathbf{I})$. This allows us to start the backward process from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively denoise for $T$ steps to obtain a sample from $\mathcal{X}$. The reverse process is parameterized as follows:
@@
Here, $p(x_{T})=\mathcal{N}(\mathbf{0},\mathbf{I}).$ Essentially, the reverse process is learnable, and $p_{\theta}(x_{t-1}\mid x_{t})$ approximates $q(x_{t-1}\mid x_{t},x_{0})$. Ho et al. (2020) show that through simple reparametrization tricks, we can convert the learning objective from approximating $q(x_{t-1}\mid x_{t},x_{0})$ to estimating the amount of noise added to go from $x_{t-1}$ to $x_{t}$. Thus, the diffusion objective is stated as minimizing the following loss function:
@@
where $t\sim\mathcal{U}(1,T)$ indicates that $t$ is sampled from a uniform distribution between 1 and $T$, $\epsilon$ is the noise added to $x_{t-1}$ to obtain $x_{t}$, and $\mathcal{\theta_{\textnormal{denoiser}}}$ is parameterized by a neural network that takes the noisy sample $x_{t}$ and the diffusion step $t$ as input to estimate $\epsilon$. This is equivalent to score-matching techniques (Song & Ermon, 2019; Song et al., 2021).
Appendix B Dataset Description
@@
In this section, we describe in detail the various datasets used in our experiments, the training, validation, and testing dataset splits, and the normalization procedure that was opted.
B.1 Electricity Dataset
@@
The electricity dataset consists of power consumption recorded for 370 users over a period of 4 years from 2011 to 2015. We frame the following task with respect to this dataset - “Generate the electricity demand pattern for the user 257, for the 3rd of August 2011,” which is a univariate time series. We consider the following features as the metadata - 370 users, 4 years, 12 months, and 31 dates (check Table 3). The power consumption is recorded every 15 minutes, so the time series is 96 timesteps long. The total number of samples without any preprocessing is 540200. We remove samples with values as 0 for the entire time series, and the resulting total number of samples is 434781. We establish a data split comprising training, validation, and test sets distributed in an 80-10-10 ratio. To obtain the split, we randomly pick 80% of the 434781 samples and assign them as the training set. The same is repeated for the validation and the test sets. We avoid using the traditional splits proposed in (Du et al., 2023) as their split creates certain year metadata features that never existed in the training set. For example, no month from 2011 exists in the training set.
B.2 Traffic Dataset
@@
For traffic volume synthesis, we use the metro interstate traffic volume dataset. The dataset has hourly traffic volume recorded from 2012 to 2018, along with metadata annotations like holidays, textual weather descriptions, weather forecasts, etc. (check Table 3). Here, we want to answer questions like - “Synthesize a traffic volume pattern for New Year’s Day, given the weather forecast”, which is a univariate time series. The dataset CSV file has a total of 48204 rows containing the traffic volume. We synthesize the traffic volume for a 96-hour window. So, to create a dataset from the CSV file, we slide a window of length 96 with a stride of 24. This gives a total of 2001 time series samples, which we randomly divide into train, validation, and test split with an 80-10-10 ratio.
B.3 Air Quality Dataset
@@
This data set contains hourly air pollutants data from 12 air-quality monitoring stations in Beijing. The meteorological data in each air-quality site are paired with the weather data from the nearest weather station (check Table 3 for more details regarding the metadata conditions). Here, the task is to synthesize a multivariate time series (6 channels) given the weather forecast metadata. The dataset has missing values, which we replace with the mean for both continuous metadata and the time series. For categorical metadata, the only missing feature is the wind direction, which we fill using an “unknown” label. The data set is split into train, validation, and test splits based on months. The recordings are available from 2013 to 2017, and we have a total of 576 months, of which we randomly pick 460 as train, 58 as validation, and 58 as test. For each month, we slide a window of length 96 with a stride of 24, and this provides a total of 12166 train time series samples, 1537 validation time series samples, and 1525 test time series samples.
B.4 ECG Dataset
@@
The PTB-XL ECG dataset is a 12-channel, 1000 time steps long, time series dataset with 17651 train, 2203 validation, and 2167 test samples. The dataset has annotated heart disease statements for each ECG time series. Here, the goal is to attempt to generate ECG time series samples for a specific heart disease statement, which is our metadata. In this work, we use 8 channels instead of 12, as shown in (Alcaraz & Strodthoff, 2023).
@@ 12 Stations, 5 years, 12 months,
Electricity 96 1 370 users, 4 years, 12 months, 31 dates NA
ECG 1000 12 71 heart disease statements NA
@@
Appendix C Metric model architecture description
@@
Figure 6: Contrastive Training of J-FTSD Feature Extractors Inspired by CLIP (Radford et al., 2021): This figure depicts the contrastive learning-based training approach for the J-FTSD feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, akin to the methodology used in CLIP. Here, we consider a time series where the first half is a triangle wave and the second half is a sine wave. The categorical metadata corresponds to this pattern, with the first half labeled as 1 (”triangle”) and the second half as 0 (”sine”). Patches of length $L_{\textnormal{patch}}$ are extracted from time series and metadata and processed through their respective feature extractors. The embeddings, $z^{c}$ from metadata and $z^{x}$ from time series, are compared using their dot products to identify correct pairings, highlighted along the matrix diagonal (in orange). The feature extractors are trained through contrastive learning, employing cross-entropy loss to enhance the accuracy of matching time series data with its relevant metadata, effectively capturing the nuanced relationship between the two.
@@
To compute our proposed J-FTSD metric, we relied on the Informer encoder architecture proposed in (Zhou et al., 2021). Specifically, we used two encoders, one for the time series and one for the metadata features, represented as $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, respectively. We made the following modifications to the Informer encoder architecture:
@@
  • The raw time series is first processed using 1D convolution layers, and we added positional encoding to the processed time series before providing as input to the self-attention layers in $\phi_{\textnormal{time}}$. We used the same positional encoding as in the Informer (Zhou et al., 2021).
@@
  • The raw metadata is processed in the same way as we processed metadata for the diffusion process, which is highlighted in Sec. 4. We individually processed or tokenized the categorical and continuous metadata using linear layers and 1D convolution layers to obtain $z$. We added positional encoding to $z$ before providing $z$ as input to the self-attention layers in $\phi_{\textnormal{meta}}$.
@@
  • We used 1D convolution layers at the end of every self-attention layer without any striding. We used striding after every 3 self-attention layers, i.e., the 1D convolution layers with stride of 2 is applied after the $3^{rd}$, $6^{th}$, $\ldots$, self-attention layers.
@@
  • At the end of the self-attention layers of both $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, we flattened the outputs and projected the outputs to a lower-dimensional space using linear layers. We used the Gaussian Error Linear Unit (GELU) activation, the same as in the Informer architecture.
@@
Now, we describe the choice of $L_{\textnormal{patch}}$ for each dataset. As explained in Sec. 5, we chose $L_{\textnormal{patch}}$ based on the minimum horizon required for a patch to contain metadata-specific features. Now, we describe the values of $L_{\textnormal{patch}}$ and the embedding size, which is the dimension of the output of the feature extractors:
@@
Table 4: Patch and Embedding sizes of all datasets.
Specifically, we chose the embedding size such that given a time series sample $x\in\mathbb{R}^{L\times F}$, where $L$ is the horizon and $F$ is the number of channels in the time series, the embedding size should be smaller than $F\times L_{\textnormal{patch}}$. This is to ensure that we are reducing the dimensionality of the time series patch.
Now, we list the hyperparameter choices, such as the number of patches from a single time series sample $N_{\textnormal{patch}}$, learning rate, etc, and the design choices in terms of the number of self-attention layers, number of transformer heads, etc.
@@
Table 5: Hyperparamters for Feature Extractors.
Appendix D Time Weaver architecture design
As mentioned in Sec. 4, Time Weaver has two denoiser backbones - CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2023). In this section, we describe the architecture changes we introduced to the CSDI and the SSSD backbones to extend their capabilities to metadata-conditioned time series generation.
D.1 Time Weaver-CSDI
Figure 7: Time Weaver-CSDI architecture: This figure shows our changes to the original conditional CSDI model (Tashiro et al., 2021). We use this model as a $\mathcal{\theta_{\textnormal{denoiser}}}$ model in our architecture, with metadata preprocessing fixed as in Fig. 3. Changes in the original architecture are colored red.
@@ D.1 Time Weaver-CSDI
  • Metadata projection: Simultaneously, we converted each categorical metadata feature in $c_{\textnormal{cat}}$ into one-hot encoding and further process using $\theta_{\textnormal{token}}^{\textnormal{cat}}$. We processed the continuous metadata, $c_{\textnormal{cont}}$, using $\theta_{\textnormal{token}}^{\textnormal{cont}}$. Both $c_{\textnormal{cat}}$ and $c_{\textnormal{cont}}$ are projected to latent representations of shape $(N_{\textnormal{batch}},L,d_{\textnormal{cat}})$ and $(N_{\textnormal{batch}},L,d_{\textnormal{cont}})$. We concatenated these latent representations along the final axis and processed them using self-attention layers, $\theta_{\textnormal{condn}}$. At the end of this preprocessing, the categorical and continuous metadata were projected to a latent representation of shape $(N_{\textnormal{batch}},L,d_{\textnormal{meta}})$. We then reshaped the projected metadata to $(N_{\textnormal{batch}},d_{\textnormal{meta}},F,L)$.
  • Diffusion step representation: The CSDI architecture represents the diffusion step using a 128-dimensional representation, which is projected to $d_{\textnormal{meta}}$. We later reshaped the diffusion step representation to $(N_{\textnormal{batch}},d_{\textnormal{meta}},F,L)$.
  • We added the input time series projection, metadata projection, and diffusion step representation and passed it through temporal and feature transformer layers in the first residual layer.
@@
For the diffusion process, our experiments with Time Weaver-CSDI use 200 diffusion steps with the noise variance schedule values of $\beta_{1}=0.0001$ and $\beta_{T}=0.1$
Now, we explain the architectural details and the corresponding hyperparameters. The number of residual layers used varies for each dataset. For the Air Quality dataset, we used 10 residual layers. Similarly, for the Traffic, Electricity, and ECG datasets, we used 8,6, and 12, respectively.
@@
D.2 Time Weaver-SSSD
Figure 8: Time Weaver-SSSD architecture: This figure shows our changes to the original conditional SSSD model (Alcaraz & Strodthoff, 2023). We use this model as a $\mathcal{\theta_{\textnormal{denoiser}}}$ model in our architecture, with metadata preprocessing being fixed as in Fig. 3. Changes to the original architecture are highlighted in red.
The Time Weaver-SSSD model is based on the structured state-space diffusion (SSSD) model (Alcaraz & Strodthoff, 2022) that was originally designed for the imputation task. The SSSD model is built on DiffWave (Kong et al., 2021) architecture. Unlike the DiffWave model, SSSD utilizes structured state-space models (SSM) (Gu et al., 2022), which connects input sequence $u(t)$ to output sequence $y(t)$ via hidden state $x(t)$. This relation can be explicitly given as:

### v2

@@ 7 Conclusion
This paper addresses a critical gap in synthetic time series data generation by introducing Time Weaver, a novel diffusion-based generative model. Time Weaver leverages heterogeneous paired metadata, encompassing categorical, continuous, and time-variant variables, to significantly improve the quality of generated time series. Moreover, we introduce a new evaluation metric, J-FTSD, to assess conditional time series generation models. This metric offers a refined approach to evaluate the specificity of generated time series relative to paired metadata conditions. Through Time Weaver, we demonstrate state-of-the-art results for generated sample quality and diversity across four diverse real-world datasets.
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Diffusion Process
Diffusion Models (DMs) are trained to denoise a noisy sample, referred to as the reverse or the backward process $p_{\theta}$. The noisy samples are generated by a Markovian forward process $q$, where we gradually corrupt a clean sample from the dataset by adding noise for $T$ diffusion steps. The forward process is predetermined by specifying a noise schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$, where $\beta_{t}\in[0,1]$. The following equations parameterize the forward process:
@@ A.1 Diffusion Process
$q(x_{t}\mid x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (8)
where $\mathcal{X}$ is the data distribution that we want to learn, $x_{0}\sim\mathcal{X}$, $\mathcal{N}(\mu,\Sigma)$ represents a Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$, and $T$ is the number of diffusion steps. The noise schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$ and $T$ are chosen such that $x_{T}$ resembles samples from a Gaussian distribution with zero mean and unit variance, i.e., $q(x_{T})\simeq\mathcal{N}(\mathbf{0},\mathbf{I})$. This allows us to start the backward process from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively denoise for $T$ steps to obtain a sample from $\mathcal{X}$. The reverse process is parameterized as follows:
@@
Here, $p(x_{T})=\mathcal{N}(\mathbf{0},\mathbf{I}).$ Essentially, the reverse process is learnable, and $p_{\theta}(x_{t-1}\mid x_{t})$ approximates $q(x_{t-1}\mid x_{t},x_{0})$. $p_{\theta}(x_{t-1}\mid x_{t})$ is parameterized using a neural network, $\mathcal{\theta_{\textnormal{denoiser}}}$. Ho et al. (2020) show that through simple reparametrization tricks, we can convert the learning objective from approximating $q(x_{t-1}\mid x_{t},x_{0})$ to estimating the amount of noise added to go from $x_{t-1}$ to $x_{t}$. Thus, the diffusion objective is stated as minimizing the following loss function:
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{x\sim\mathcal{X},\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I}),t\sim\mathcal{U}(1,T)}\left[\|\epsilon-\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t)\right\|^{2}_{2}],$ (10)
@@
where $t\sim\mathcal{U}(1,T)$ indicates that $t$ is sampled from a uniform distribution between 1 and $T$, $\epsilon$ is the noise added to $x_{t-1}$ to obtain $x_{t}$, and $\mathcal{\theta_{\textnormal{denoiser}}}$ takes the noisy sample $x_{t}$ and the diffusion step $t$ as input to estimate $\epsilon$. This is equivalent to the score-matching techniques (Song & Ermon, 2019; Song et al., 2021).
A.2 Dataset Description
@@
In this section, we describe in detail the various datasets used in our experiments, their training, validation, and testing splits, and the normalization procedures.
A.2.1 Electricity Dataset
@@
The Electricity dataset consists of power consumption recordings for 370 users over four years from 2011 to 2015. We frame the following task for this dataset - “Generate the electricity demand pattern for user 257, for the 3rd of August 2011,” which is a univariate time series. We consider the following features as the metadata: 370 users, four years, 12 months, and 31 dates (Table 3). The power consumption is recorded every 15 minutes, so the time series is 96 time steps long. The total number of samples without any preprocessing is 540200. We remove samples with values of 0 for the entire time series, and the resulting total number of samples is 434781. We establish a data split comprising training, validation, and test sets distributed in an 80-10-10 ratio. To obtain the split, we randomly pick 80% of the 434781 samples and assign them to the training set. The same is repeated for the validation and the test sets. We avoid using the traditional splits proposed by Du et al. (2023), as their split creates certain year metadata features in the test set that do not exist in the training set. For example, no month from 2011 exists in the training set split proposed by Du et al. (2023).
A.2.2 Traffic Dataset
@@
For traffic volume synthesis, we use the metro interstate traffic volume dataset. The dataset has hourly traffic volume recordings from 2012 to 2018, along with metadata annotations like holidays, textual weather descriptions, and weather forecasts (Table 3). Here, we want to generate time series samples for prompts such as “Given the following weather forecast, synthesize a traffic volume pattern for New Year’s Day.”, which is a univariate time series. The dataset CSV file has 48204 rows containing the traffic volume. We synthesize the traffic volume for a 96-hour window. So, to create a dataset from the CSV file, we slide a window of length 96 with a stride of 24. This gives a total of 2001 time series samples, which we randomly divide into train, validation, and test sets with an 80-10-10 ratio.
A.2.3 Air Quality Dataset
@@
This dataset contains hourly air pollutants data from 12 air quality monitoring stations in Beijing. The meteorological data in each air quality site are paired with the weather data from the nearest weather station (please refer to Table 3 for more details regarding the metadata conditions). Here, the task is to synthesize a multivariate (6 channels) time series given the weather forecast metadata. The dataset has missing values, which we replace with the mean for both continuous metadata and the time series. For categorical metadata, missing values exist only in the wind direction metadata feature, which we fill using an “unknown” label. The dataset is split into train, validation, and test sets based on months. The recordings are available from 2013 to 2017, and we have a total of 576 months, of which we randomly picked 460 as train data, 58 as validation data, and 58 as test data. For each month, we slide a window of length 96 with a stride of 24, and this provides a total of 12166 train time series samples, 1537 validation time series samples, and 1525 test time series samples.
A.2.4 ECG Dataset
@@
The PTB-XL ECG dataset is a 12-channel (1000 time steps long) time series dataset with 17651 train, 2203 validation, and 2167 test samples. The dataset has annotated heart disease statements for each ECG time series. Here, the goal is to generate ECG time series samples for a specific heart disease statement, which is our metadata. In this work, we use 8 channels instead of 12, as shown in (Alcaraz & Strodthoff, 2023).
@@ 12 Stations, 5 years, 12 months,
Electricity 96 1 370 users, 4 years, 12 months, 31 dates N.A.
ECG 1000 8 71 heart disease statements N.A.
@@
A.3 Architecture Description Of The Feature Extractors Used In J-FTSD
@@
Figure 6: Contrastive Training of J-FTSD Feature Extractors Inspired By CLIP (Radford et al., 2021): This figure depicts the contrastive learning-based training approach for the J-FTSD feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, akin to the methodology used in CLIP. Here, we consider a time series where the first half is a triangle wave and the second half is a sine wave. The categorical metadata corresponding to this time series has the first half labeled as 1 (“triangle”) and the second half as 0 (“sine”). Patches of length $L_{\textnormal{patch}}$ are extracted from time series and metadata and processed through their respective feature extractors. The embeddings, $z_{c}$ from the metadata encoder and $z_{x}$ from the time series encoder, are compared using their dot products to identify correct pairings, highlighted along the matrix diagonal (in orange). The feature extractors are trained through contrastive learning, employing cross-entropy loss to enhance the accuracy of matching time series data with its relevant metadata, effectively capturing the nuanced relationship between the two.
@@
To compute our proposed J-FTSD metric, we relied on the Informer encoder architecture proposed by Zhou et al. (2021). Specifically, we used two encoders, one for the time series and one for the metadata features, represented as $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, respectively. We made the following modifications to the Informer encoder architecture:
@@
  • The raw time series was first processed using 1D convolution layers. We added positional encoding to the processed time series before providing it as input to the self-attention layers in $\phi_{\textnormal{time}}$. We used the same positional encoding as suggested by Zhou et al. (2021).
@@
  • The raw metadata was processed in the same way as the metadata for the time series generation process, which is highlighted in Sec. 4. We individually processed or tokenized the categorical and continuous metadata using linear layers and 1D convolution layers to obtain the metadata embedding $z$. We added positional encoding to $z$ before providing $z$ as input to the self-attention layers in $\phi_{\textnormal{meta}}$.
@@
  • We used 1D convolution layers at the end of every self-attention layer. We used striding after every three self-attention layers, i.e., 1D convolution layers with a stride of 2 were applied after the $3^{rd}$, $6^{th}$, $\ldots$, self-attention layers.
@@
  • At the end of the self-attention layers of both $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, we flattened the outputs and projected the outputs to a lower-dimensional space using linear layers. We used the Gaussian Error Linear Unit (GELU) activation, similar to the Informer architecture.
@@
Now, we describe the choice of $L_{\textnormal{patch}}$ for each dataset. As explained in Sec. 5, we chose $L_{\textnormal{patch}}$ based on the minimum horizon required for a patch to contain metadata-specific features. For each dataset, the values of $L_{\textnormal{patch}}$ and the embedding size (the output dimension of the feature extractors) are tabulated in Table 4.
@@
Table 4: Choice of the patch and embedding sizes. In this table, we list the choices of the patch size $L_{\textnormal{patch}}$ and the embedding size used for each dataset in our experiments.
Specifically, we chose the embedding size such that given a time series sample $x\in\mathbb{R}^{L\times F}$, where $L$ is the horizon, and $F$ is the number of channels in the time series, the embedding size should be smaller than $F\times L_{\textnormal{patch}}$. This is to ensure that we are reducing the dimensionality of the time series patch.
Now, we list the hyperparameter choices used for training the feature extractors in Table 5. These include the number of patches from a single time series sample $N_{\textnormal{patch}}$, learning rate, etc, and the design choices in terms of the number of self-attention layers, number of transformer heads, etc.
Fig. 6 provides a pictorial representation of the contrastive learning framework used to train the feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. The training process ensures that the time series samples with similar paired metadata have similar projections in the latent space.
@@
Table 5: Hyperparameters for the feature extractors.
A.4 Time Weaver Architecture Design
As mentioned in Sec. 4, Time Weaver has two denoiser backbones - CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2023). In this section, we describe the architecture changes we introduced to the CSDI and the SSSD backbones to extend their capabilities to metadata-specific time series generation.
A.4.1 Time Weaver-CSDI
Figure 7: Time Weaver-CSDI architecture: This figure shows our changes to the original conditional CSDI model (Tashiro et al., 2021). We use this model as our denoiser ($\mathcal{\theta_{\textnormal{denoiser}}}$) in our architecture, with metadata preprocessing fixed as in Fig. 3. Changes to the original architecture are colored in red.
@@ A.4.1 Time Weaver-CSDI
  • Metadata projection: Simultaneously, we converted each categorical metadata feature in $c_{\textnormal{cat}}$ into one-hot encoding and further processed using $\theta_{\textnormal{token}}^{\textnormal{cat}}$. Similarly, we processed the continuous metadata, $c_{\textnormal{cont}}$, using $\theta_{\textnormal{token}}^{\textnormal{cont}}$. Both $c_{\textnormal{cat}}$ and $c_{\textnormal{cont}}$ were projected to latent representations of shape $(N_{\textnormal{batch}},L,d_{\textnormal{cat}})$ and $(N_{\textnormal{batch}},L,d_{\textnormal{cont}})$, respectively. These latent representations were concatenated along the final axis and processed using the self-attention layer $\theta_{\textnormal{condn}}$. At the end of this preprocessing, the categorical and continuous metadata were projected to a latent representation of shape $(N_{\textnormal{batch}},L,d_{\textnormal{meta}})$. We then broadcasted and reshaped the projected metadata to $(N_{\textnormal{batch}},d_{\textnormal{meta}},F,L)$.
  • Diffusion step representation: The CSDI architecture represents the diffusion step using a 128-dimensional representation, which is projected to $d_{\textnormal{meta}}$. We broadcasted and reshaped the diffusion step representation to $(N_{\textnormal{batch}},d_{\textnormal{meta}},F,L)$.
  • Then, we added the input time series projection, metadata projection, and diffusion step representation. This sum was passed to the temporal and feature transformer layers in the first residual layer.
@@
These modifications are highlighted in red in Fig. 7. For the diffusion process, our experiments with Time Weaver-CSDI used 200 diffusion steps with the noise variance schedule values of $\beta_{1}=0.0001$ and $\beta_{T}=0.1$.
Now, we list the architectural choices and the corresponding hyperparameter choices in Table 6. The number of residual layers used varies for each dataset. For the Air Quality dataset, we used 10 residual layers. Similarly, for the Traffic, Electricity, and ECG datasets, we used 8, 6, and 10 residual layers, respectively.
@@
A.4.2 Time Weaver-SSSD
Figure 8: Time Weaver-SSSD architecture: This figure shows our changes to the original conditional SSSD model (Alcaraz & Strodthoff, 2023). We use this model as the denoiser ($\mathcal{\theta_{\textnormal{denoiser}}}$) in our architecture, with metadata preprocessing being fixed as in Fig. 3. Changes to the original architecture are highlighted in red.
The Time Weaver-SSSD model is based on the structured state-space diffusion (SSSD) model (Alcaraz & Strodthoff, 2022) which was originally designed for imputation tasks. The SSSD model is built on the DiffWave (Kong et al., 2021) architecture. Unlike the DiffWave model, SSSD utilizes the structured state-space model (SSM) (Gu et al., 2022), which connects the input sequence $u(t)$ to the output sequence $y(t)$ via the hidden state $x(t)$. This relation can be explicitly given as: